OpenStudy (anonymous):

Solve for x 3^x = e^2

OpenStudy (anonymous):

\[3^{x} = e ^{2}\] Show work

OpenStudy (lgbasallote):

i would get the square root of both sides first to make e alone... \[\large \sqrt{3^x} = e\] do you understand that?

OpenStudy (anonymous):

I don't have to use ln?

OpenStudy (lgbasallote):

you will

OpenStudy (lgbasallote):

i'll simplify this into \[\large 3^{x/2} = e\] you understand that algebra right?

OpenStudy (anonymous):

first of all e=2.71828 so 2.71828^2=3^x 7.38905=3^x x=1.82048

OpenStudy (anonymous):

Where did you get the 7.389..? When do I use ln?

OpenStudy (lgbasallote):

\[\large 3^{x/2} = e\] \[\log_e 3 = \frac{x}{2}\] \[\ln 3 = \frac{x}{2}\] \[x = 2\ln 3\] \[x = \ln 3^2\] \[x = \ln 9\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3^x+%3D+e^2

OpenStudy (anonymous):

Its telling me the answer should be 2/ln 3= 1.820

OpenStudy (anonymous):

take the natural log to start off with right away. you'll get ln3^x = 2 when you have a log raised to a power, you can bring the power down x * ln3 = 2 get x by itself x = 2/ln3 ta - daaaaaaaaaaa

OpenStudy (anonymous):

:D lol thanks!